OpenStudy (anonymous):

x/4+12=5 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

OpenStudy (luigi0210):

\[\frac{ x }{ 4 }+12=5\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

I think its 68

OpenStudy (jhannybean):

\[\large \cfrac{x}{4} + 12 = 5\]\[\large 4\left(\cfrac{x}{4}+12 = 5\right)\]\[\large x+ 48 =20\]\[\large x= 20-48\]

OpenStudy (jhannybean):

Elaborating on what @Luigi0210 was working with.

OpenStudy (anonymous):